Jo invested part of her $8000 at 12% and part at 8%. If the annual investment income is $825, how much was invested at each rate?
Answer by mananth(16946)

Investment I 12.00% per annum ------------- Amount invested =x
Investment II 8.00% per annum ------------ Amount invested = y
8000
Interest----- 825

Investment I 12.00% per annum ---x
Investment II 8.00% per annum ---y
Credit on both cards
x + 1 y= 8000 -------------1
Interestto be paid on both
12.00% x + 8.00% y= 825
Multiply by 100
12 x + 8 y= 82500.00 --------2
Multiply (1) by -12
we get
-12 x -12 y= -96000.00
Add this to (2)
0 x -4 y= -13500
divide by -4
y = 3375
Investment I 12.00% $ 4625
Investment II 8.00% $ 3375

CHECK
4625 --------- 12.00% ------- 555.00
3375 ------------- 8.00% ------- 270.00
Total -------------------- 825.00

m.ananth@hotmail.ca